OpenStudy (anonymous):

give exact and approximate solutions to three decimal place (x-2)^2=64

OpenStudy (anonymous):

(x-2)^2=64 (x-2)=+ or - 8 x=-6 and 10

OpenStudy (anonymous):

so = -6.000 and 10.000

OpenStudy (mertsj):

Taking the square root of both sides yields x - 2 = 8 or -8. Add 2 to both sides. x = 10 or -6. Didn't you learn anything from the previous example?

OpenStudy (anonymous):

lol obviously not otherwise he wouldnt ask it haha he just wants the answers